Parramore Corp has $13 million of sales, $3 million of inventories, $4 million of receivables, and $2 million of payables. Its cost of goods sold is 75% of sales, and it finances working capital with bank loans at an 8% rate. Assume 365 days in year for your calculations.

  1. What is Parramore's cash conversion cycle (CCC)? Do not round intermediate calculations. Round your answer to two decimal places.
      days

  2. If Parramore could lower its inventories and receivables by 11% each and increase its payables by 11%, all without affecting sales or cost of goods sold, what would be the new CCC? Do not round intermediate calculations. Round your answer to two decimal places.
      days

  3. How much cash would be freed up, if Parramore could lower its inventories and receivables by 11% each and increase its payables by 11%, all without affecting sales or cost of goods sold? Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000. Do not round intermediate calculations. Round your answer to the nearest dollar.
    $   

  4. By how much would pretax profits change, if Parramore could lower its inventories and receivables by 11% each and increase its payables by 11%, all without affecting sales or cost of goods sold? Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000. Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

Answer 1: Cash conversion cycle = DIO + DSO – DPO

DIO = Days Inventory Outstanding = (Average Inventory / Cost of goods sold)*365 = $3 Million / ($13 Million *75%) *365 = 112.3077 days

DSO = Days Sales Outstanding = (Average Accounts Receivables / Total Credit sales)*365 = ($4 Million / $ 13 Million) *365 = 112. 3077 days

DPO = Days Payable Outstanding = (Average Accounts Payable / Cost of goods sold)*365 = $ 2 Million / ($13 Million *75%) *365 = 74.8718 days

Cash conversion cycle = 112.3077+112.3077-74.8718 = 149.74 Days

Answer 2: If inventories and receivables lowered and accounts payable increase by 11% then new figures will be:

Inventories = $ 3 Million *(1-.11) = 2.67 Million

Accounts receivables = $ 4 Million * (1-0.11) = 3.56 Million

Accounts payable = $ 2 Millions * (1+0.11) = 2.22 Million

Sales = $13 Million (same)

COGS = $ 13 Million *75% = 9.75 Million (same)

DIO = 2.67 Million / 9.75 Million = 99.95384615 Days

DSO = 3.56 Million / 13 Million = 99.95384615 Days

DPO = 2.22 Million / 9.75 Million = 83.10769231 Days

NEW cash conversion cycle = 99.95384615+99.95384615-83.10769231 = 116.80 Days

Answer 3:

Cash /investment being ideal in normal position = 3+4-2 = 5 Million

Cash /investment being ideal in new scenario = 2.67+3.56-2.22 = 4.01Million

So cash freed = 5 Million – 4.01 Million = 0.99 Million = 990,000

Let understand this further

Particulars

Old Position

New Position

Cash freed

Cash Converted from inventories

3

2.67

0.33

Cash Collected from Receivables

4

3.56

0.44

Cash not paid to suppliers

2

2.22

0.22

0.99

Answer 4:

Impact on profit:

Since we know that 990,000 cash will be freed if the given situation is implied, so there will be saving of interest @ 8% on 990,000

Reduction in interest cost = Increase in profit = 990,000*8% = 79,200